Carbon tetrachloride can be formed by reacting chlorine with methane, calculate the enthalpy of reaction.

CH4+2Cl2-->CCl4+2H2


1
Expert's answer
2021-05-18T07:48:51-0400

To find the enthalpy of the reaction, the values of standard enthalpies of formation of all participating compounds should be given. They are not provided here, but can be found in numerous open sources. Also, since the phases are not mentioned, we assume that the compounds are in their standard state (CCl4 liquid, and the rest - gases).

"\Delta{H_f}(CCl_4(l))=-128.4kJ/mol"

"\Delta{H_f}(CH_4(g))=-74.8kJ/mol"

"\Delta{H_f}" for Cl2 and H2 are in fact zero (for all compounds made of one type of atoms).


"\Delta{H_{reaction}}=\sum{\Delta{H_f(products)-\sum{\Delta{H_f}(reactants)}}}=\Delta{H_f(CCl_4)-\Delta{H_f}(CH_4)}=-128.4-(-74.8)=-53.6kJ/mol"


Answer: -53.6 kJ/mol